# Kermut: Composite kernel regression for protein variant effects

### v2

@@ B.3.1 Example wall-clock time
Appendix C Data
@@ Appendix E Detailed results
The ProteinGym suite provides an aggregation procedure, whereby the predictive performance across both cross-validation schemes and functional categories can be gauged. We provide these results in Tables 5, 6, 7 and 8.
Table 5: Aggregated Spearman results on the ProteinGym substitution benchmark. Performance is shown per cross-validation scheme. Kermut reaches superior performance across the board. The fifth data column shows the non-parametric bootstrap standard error of the difference between the Spearman performance for each model and Kermut, computed over 10,000 bootstrap samples from the set of proteins in the ProteinGym substitution benchmark.
@@
Table 6: Aggregated Spearman results on the ProteinGym substitution benchmark. Performance is shown per functional category. Kermut reaches superior performance across the board.
@@
Table 7: Aggregated MSE results on the ProteinGym substitution benchmark. Performance is shown per cross-validation scheme. Kermut reaches superior performance across the board. The fifth data column shows the non-parametric bootstrap standard error of the difference between the MSE performance for each model and Kermut, computed over 10,000 bootstrap samples from the set of proteins in the ProteinGym substitution benchmark.
@@
Tranception Emb. 0.972 0.833 0.503 0.769 0.024
ESM-1v Emb. 0.937 0.861 0.563 0.787 0.031
TranceptEVE + OHE 0.953 0.914 0.743 0.870 0.020
MSAT + OHE 0.963 0.934 0.749 0.882 0.021
DeepSequence + OHE 0.967 0.940 0.767 0.891 0.018
Table 8: Aggregated results on the ProteinGym substitution benchmark. Performance is shown per functional category. Kermut reaches superior performance across the board.
@@ Appendix F Ablation results
In Section 4.1, an ablation study was carried out by removing components of Kermut. In Table 2, the performance difference in Spearman correlation was shown. In Table 9 we see the performance difference in MSE. The absolute Spearman and MSE values are shown in Tables 10 and 11.
Table 9: Ablation results. Key components of the kernel are removed and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the GP formulation. The metrics are subtracted from Kermut to show the change in performance. Positive $\Delta$MSE values indicate drop in performance.
@@
Table 10: Ablation results. Key components of the kernel are removed and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the kernel formulation
Description Ablation Spearman ($\uparrow$)
@@
No structure kernel $k_{\text{struct}}=0$ $0.523$ $0.550$ $0.710$ $0.594$
No sequence kernel $k_{\text{seq}}=0$ $0.567$ $0.581$ $0.695$ $0.614$
No inter-residue dist. $k_{d}=1$ $0.557$ $0.577$ $0.740$ $0.625$
No mut. prob./site comp. $k_{p}=k_{H}=1$ $0.570$ $0.590$ $0.736$ $0.632$
Const. mean $m(\mathbf{x})=\alpha$ $0.568$ $0.595$ $0.735$ $0.633$
No mut. prob. $k_{p}=1$ $0.584$ $0.609$ $0.737$ $0.643$
No site comp. $k_{H}=1$ $0.600$ $0.619$ $0.743$ $0.654$
Kermut $0.607$ $0.628$ $0.747$ $0.660$
Table 11: Ablation results. Key components of the kernel are removed and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the kernel formulation
Description Ablation MSE ($\downarrow$)
@@
No structure kernel $k_{\text{struct}}=0$ $0.825$ $0.769$ $0.460$ $0.685$
No sequence kernel $k_{\text{seq}}=0$ $0.789$ $0.741$ $0.491$ $0.674$
No inter-residue dist. $k_{d}=1$ $0.787$ $0.742$ $0.425$ $0.652$
No mut. prob./site comp. $k_{p}=k_{H}=1$ $0.775$ $0.722$ $0.438$ $0.645$
Const. mean $m(\mathbf{x})=\alpha$ $0.771$ $0.718$ $0.429$ $0.639$
No mut. prob. $k_{p}=1$ $0.761$ $0.703$ $0.436$ $0.633$
No site comp. $k_{H}=1$ $0.734$ $0.691$ $0.422$ $0.615$
Kermut $0.729$ $0.682$ $0.414$ $0.609$
Appendix G Results for multi-mutants
@@
Table 12: Results in multi-mutant setting. Each row corresponds to a different setting of training and evaluation domain. Third row corresponds to the fold_rand_multiples split-scheme from ProteinGym. Experiments are carried out on 51 datasets, corresponding to all datasets with multiple mutants with less than 7500 variants in total with the exception of GCN4_YEAST_Staller_2018, which has been removed due to its high mutation count. : Constant mean.
@@
Domain $k_{\text{seq}}$ Kermut Kermut $k_{\text{seq}}$ Kermut Kermut
@@
Appendix H Histogram over assays for multi-mutant datasets
A histogram over assay values for 51/69 dataset with multi-mutants can be seen in Figure 4. The histograms are colored according to the number of mutations per variant. The assay distribution belong to different modalities depending on the number of mutations present, where double mutations often lead to a loss of fitness.
Figure 4: Histogram over assay values for 51/69 datasets with multi-mutants. All datasets with more than 7500 variants are ignored. The dataset GCN4_YEAST_Staller_2018 is removed due to its high mutation count which skews the figure. The histograms are colored according to the number of mutations per variant. The assay distribution belong to different modalities depending on the number of mutations present, where double mutations often lead to a loss of fitness.
Appendix I Uncertainty calibration
I.1 Confidence interval-based calibration
@@
I.2 Error-based calibration
@@
I.3 Uncertainty calibration across all datasets
@@
For each dataset, split, and fold, we perform a linear regression to the error-based calibration curves and summarize the slope and intercept. As perfect calibration corresponds to a diagonal line, we want the distribution over the slopes to be centred on one and the distribution over intercepts to be centred on zero. Boxplots for this analysis can be seen in Figure 6. As indicated in (a), most of Kermut’s calibration curves are well-behaved, while the baseline GP in (b) is generally poorly calibrated.
Figure 5: Calibration metrics per domain for Kermut and the sequence kernel on ESM-2 embeddings. Random, modulo, and contiguous domains are from the ProteinGym substitution benchmark. Multiples corresponds to training and testing on both single and double mutants. Extrapolation corresponds to training on singles and predicting on doubles. 51 datasets with multi-mutants was used for the figure for all domains for comparability. The performance results for the multi-mutant setting can be found in Table 12. Errorbars correspond to standard error.
@@
Figure 6: Boxplot of intercepts and slopes of error-based calibration curves for Kermut and a baseline GP with the sequence kernel on ESM-2 embeddings. Perfect calibration has an intercept of zero and a slope of one (indicated by dashed lines). The baseline GP has poor calibration compared to Kermut. Horizontal lines indicate 0.9, 0.75, 0.5, 0.25, 0.1 quantiles.
Appendix J Predicted vs. true values for calibration datasets
J.1 BLAT_ECOLX_Stiffler_2015
Figure 7: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
J.2 PA_I34A1_Wu_2015
Figure 8: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
J.3 TCRG1_MOUSE_Tsuboyama_2023
Figure 9: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
J.4 OPSD_HUMAN_Wan_2019
Figure 10: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
Appendix K Calibration curves for ProteinNPT
K.1 ProteinNPT details
@@
Figure 11: Calibration curves for ProteinNPT on the four dataset from Table 3. Standard deviation over CV folds is shown as vertical bars. Perfect calibration corresponds to diagonal lines ($y=x$) and is shown as dashed lines in each plot. The predictive uncertainties for ProteinNPT are very small, resulting in poor out-of-the-box calibration as seen on the x-axis in (b) and in Figures 12, 13, 15 and 14. However, as indicated by the trend in both (a) and (b), the errors correlate with the magnitude of the uncertainties. This suggests that a recalibration might be sufficient to achieve good calibration.
K.3 Predicted vs. true values for calibration datasets
K.3.1 BLAT_ECOLX_Stiffler_2015
Figure 12: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$). While the predictions are good, the model is very overconfident.
K.3.2 PA_I34A1_Wu_2015
Figure 13: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$). Despite the relatively poor predictions, the model remains overconfident.
K.3.3 TCRG1_MOUSE_Tsuboyama_2023
Figure 14: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$). While the predictions are good, the model is very overconfident. Despite the relatively poor predictions, the model remains overconfident.
K.3.4 OPSD_HUMAN_Wan_2019
Figure 15: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
Appendix L Alternative zero-shot methods
@@
Table 13: Performance using alternate zero-shot methods. The experiments are carried out on 174/217 datasets. : ESM-2 in this table is equivalent to Kermut from the main results in Table 1.
@@
EVE 0.610 0.628 0.750 0.663 0.729 0.680 0.412 0.607
ESM-2 0.607 0.628 0.747 0.660 0.729 0.682 0.414 0.609
VESPA 0.605 0.622 0.742 0.657 0.736 0.697 0.424 0.619
GEMME 0.604 0.622 0.744 0.657 0.729 0.682 0.416 0.609
TranceptEVE L 0.597 0.617 0.744 0.653 0.742 0.694 0.420 0.618
ESM-IF1 0.581 0.605 0.738 0.641 0.757 0.708 0.424 0.630
ProteinMPNN 0.573 0.599 0.734 0.635 0.769 0.717 0.429 0.638
Constant mean 0.568 0.595 0.735 0.633 0.771 0.718 0.429 0.639
Appendix M Ethics
We have introduced a general framework to predict variant effects given labelled data. The intent of our work is to use the framework to model and subsequently optimize proteins. We acknowledge that – in principle – any protein property can be modelled (depending on the available data), which means that potentially harmful proteins can be engineered using our method. We encourage the community to use our proposed method for beneficial purposes only, such as the engineering of efficient enzymes or for the characterization of potentially pathogenic variants for the betterment of biological interpretation and clinical treatment.
Appendix N License
The codebase will be made publicly available under the open source MIT License once anonymity is not an issue.
@@
We show the performance of Kermut and ProteinNPT per DMS assay in Figures 16, 17, 18 and 19. The figures show the average performance and the performance per split, respectively.
@@
Figure 16: Spearman’s correlation per DMS assay, averaged over the three split schemes. The figure has been split in two to fit.
@@
Figure 17: Spearman’s correlation per DMS assay in the random split scheme. The figure has been split in two to fit. The performance difference between Kermut and ProteinNPT for the random split is smallest (a), i.e., for the assays where the performance is high.

### v3

@@ C.4.1 Example wall-clock time
C.5 Computational complexity
Evaluating the kernel for two variants $\mathbf{x}$ and $\mathbf{x}^{\prime}$, involves computation of the two components $k_{\text{seq}}$ and $k_{\text{struct}}$:
For each variant $k_{\text{seq}}$ requires a forward pass through ESM-2, which is based on the transformer architecture and has quadratic scaling with respect to the sequence length. Given the ESM-2 embeddings, the computational complexity is constant in sequence length and number of mutations.
$k_{\text{struct}}$ requires a single forward pass through ProteinMPNN for the wild type protein. Given the output of ProteinPMNN, the computational complexity for evaluation of the kernel is $m_{1}\times m_{2}$ for two proteins with $m_{1}$ and $m_{2}$ number of mutations. The computational complexity of $k_{\text{struct}}^{1}$ is constant in sequence length and number of mutations.
Appendix D Data
@@ Appendix E Detailed results
The ProteinGym suite provides an aggregation procedure, whereby the predictive performance across both cross-validation schemes and functional categories can be gauged. We provide these results in Tables E.1, E.2, E.3 and E.4. We mirror the label normalization from [19] and [7], where, for each fold of cross-validation, train and test labels are normalized given the mean and standard deviation of the training labels. The Spearman correlation coefficient and MSE are then computed in the normalized space across folds, leading to a single Spearman coefficient and MSE per assay per split (i.e., not as an average of metrics per CV-fold).
Table E.1: Aggregated Spearman results on the ProteinGym substitution benchmark. Performance is shown per cross-validation scheme. Kermut reaches superior performance across the board. The fifth data column shows the non-parametric bootstrap standard error of the difference between the Spearman performance for each model and Kermut, computed over 10,000 bootstrap samples from the set of proteins in the ProteinGym substitution benchmark.
@@
Table E.2: Aggregated Spearman results on the ProteinGym substitution benchmark. Performance is shown per functional category. Kermut reaches superior performance across the board.
@@
Table E.3: Aggregated MSE results on the ProteinGym substitution benchmark. Performance is shown per cross-validation scheme. Kermut reaches superior performance across the board. The fifth data column shows the non-parametric bootstrap standard error of the difference between the MSE performance for each model and Kermut, computed over 10,000 bootstrap samples from the set of proteins in the ProteinGym substitution benchmark.
@@
Tranception Emb. 0.972 0.833 0.503 0.769 0.023
ESM-1v Emb. 0.937 0.861 0.563 0.787 0.030
TranceptEVE + OHE 0.953 0.914 0.743 0.870 0.019
MSAT + OHE 0.963 0.934 0.749 0.882 0.020
DeepSequence + OHE 0.967 0.940 0.767 0.891 0.017
Tranception + OHE 0.985 0.934 0.766 0.895 0.022
ESM-1v + OHE 0.977 0.949 0.764 0.897 0.013
OHE 1.158 1.125 0.898 1.061 0.017
Table E.4: Aggregated results on the ProteinGym substitution benchmark. Performance is shown per functional category. Kermut reaches superior performance across the board.
@@
Tranception + OHE 0.831 1.246 0.787 0.845 0.765
ESM-1v + OHE 0.843 1.192 0.795 0.870 0.783
OHE 1.022 1.306 0.986 1.040 0.949
E.1 Results on new splits
The modulo and contiguous splits in ProteinGym were updated in April 2024. For completeness, we here provide results using Kermut on the updated splits for all 217 DMS assays. These can be seen in Table E.5, where we observe a slight decrease in performance for the contiguous and modulo splits, compared to the reference results in Table 1.
Table E.5: Performance on the ProteinGym benchmark using the corrected splits.
Model Model name Spearman ($\uparrow$) MSE ($\downarrow$)
type Contig. Mod. Rand. Avg. Contig. Mod. Rand. Avg.
GP Kermut 0.591 0.631 0.744 0.655 0.739 0.680 0.141 0.611
@@ Appendix F Ablation results
In Section 4.1, an ablation study was carried out by removing components of Kermut. In Table 2, the performance difference in Spearman correlation was shown. In Table F.1 we see the performance difference in MSE. The aggregated absolute Spearman and MSE values are shown in Tables F.2 and F.3, while we show the performance per functional category in Tables F.4 and F.5. These results suggest that an even wider combinatorial examination of Kermut’s kernel composition might lead to slightly increased performance, e.g., by multiplying a Matérn 5/2 sequence kernel with the structure kernel. We must however note that the standard errors in Tables F.2 and F.3 suggest that while the product and Matérn configurations lead to better results on the ablation datasets, the differences are not significant.
In addition to the shown methods, we considered sequence-only kernels as a baselines. One example is the inverse-multiquadratic Hamming (IMQ-H) kernel from [31]. This kernel, however, relies on the Hamming distance between one-hot encoded sequences for its covariances. For ProteinGym’s single-mutant benchmark, this is not sufficient as the Hamming distance between all variants is 2, resulting in constant predictions for all folds and subsequent Spearman correlations narrowly centered on 0.
Table F.1: Ablation results. Key components of the kernel are removed and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the GP formulation. The metrics are subtracted from Kermut to show the change in performance. Positive $\Delta$MSE values indicate drop in performance.
@@
Table F.2: Ablation results. Key components of the kernel are removed or altered and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the kernel formulation.
Description Ablation Spearman ($\uparrow$) Std. err.
@@
No structure kernel $k_{\text{struct}}=0$ 0.523 0.550 0.710 0.594 0.009
No sequence kernel $k_{\text{seq}}=0$ 0.565 0.582 0.695 0.614 0.006
No inter-residue dist. $k_{d}=1$ 0.556 0.577 0.739 0.624 0.004
No mut. prob./site comp. $k_{p}=k_{H}=1$ 0.570 0.591 0.737 0.633 0.006
Const. mean $m(\mathbf{x})=\alpha$ 0.569 0.596 0.735 0.633 0.007
No mut. prob. $k_{p}=1$ 0.583 0.609 0.738 0.643 0.004
No site comp. $k_{H}=1$ 0.599 0.619 0.743 0.653 0.002
Kermut (SE in $k_{H}$) $k_{H}=k_{\text{SE}}$ 0.598 0.621 0.745 0.655 0.002
Kermut (JSD in $k_{H}$) $k_{H}=k_{\text{JSD}}$ 0.604 0.626 0.745 0.658 0.001
[…]
Kermut (product) $k=k_{\text{struct}}\times k_{\text{seq}}$ 0.606 0.632 0.750 0.662 0.003
Kermut 0.605 0.628 0.743 0.659 0.000
Table F.3: Ablation results. Key components of the kernel are removed or altered and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the kernel formulation.
Description Ablation MSE ($\downarrow$) Std. err.
@@
No structure kernel $k_{\text{struct}}=0$ 0.825 0.769 0.460 0.684 0.012
No sequence kernel $k_{\text{seq}}=0$ 0.791 0.744 0.492 0.676 0.008
No inter-residue dist. $k_{d}=1$ 0.789 0.743 0.426 0.652 0.006
No mut. prob./site comp. $k_{p}=k_{H}=1$ 0.775 0.722 0.436 0.644 0.007
Const. mean $m(\mathbf{x})=\alpha$ 0.770 0.718 0.429 0.639 0.005
No mut. prob. $k_{p}=1$ 0.761 0.704 0.436 0.634 0.006
No site comp. $k_{H}=1$ 0.735 0.691 0.421 0.616 0.002
Kermut (SE in $k_{H}$) $k_{H}=k_{\text{SE}}$ 0.738 0.690 0.416 0.614 0.003
Kermut (JSD in $k_{H}$) $k_{H}=k_{\text{JSD}}$ 0.731 0.684 0.415 0.610 0.002
@@
Kermut (product) $k=k_{\text{struct}}\times k_{\text{seq}}$ 0.726 0.678 0.411 0.605 0.004
Kermut 0.730 0.683 0.420 0.611 0.000
Table F.4: Ablation results. Key components of the kernel are removed or altered and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the kernel formulation. Performance is shown per functional category.
Model name Ablation Spearman per function ($\uparrow$)
Activity Binding Expression Fitness Stability
Const. mean $m(\mathbf{x})=\alpha$ 0.579 0.580 0.651 0.536 0.821
No site comp. $k_{H}=1$ 0.590 0.619 0.664 0.574 0.820
No mut. prob. $k_{p}=1$ 0.590 0.611 0.651 0.564 0.801
No mut. prob./site comp. $k_{p}=k_{H}=1$ 0.569 0.601 0.648 0.557 0.790
No inter-residue dist. $k_{d}=1$ 0.562 0.560 0.634 0.546 0.818
No sequence kernel $k_{\text{seq}}=0$ 0.578 0.616 0.611 0.547 0.716
No structure kernel $k_{\text{struct}}=0$ 0.531 0.529 0.614 0.519 0.778
[…]
Kermut (SE in $k_{H}$) $k_{H}=k_{\text{SE}}$ 0.593 0.623 0.664 0.576 0.818
Kermut (JSD in $k_{H}$) $k_{H}=k_{\text{JSD}}$ 0.599 0.628 0.664 0.578 0.823
Kermut (product) $k=k_{\text{struct}}\times k_{\text{seq}}$ 0.599 0.632 0.674 0.578 0.829
Kermut 0.602 0.625 0.665 0.578 0.824
Table F.5: Ablation results. Key components of the kernel are removed or altered and the model is trained and evaluated on 174/217 assays from the ProteinGym benchmark. The ablation column shows the alteration to the kernel formulation. Performance is shown per functional category.
Model name Ablation MSE per function ($\downarrow$)
Activity Binding Expression Fitness Stability
Const. mean $m(\mathbf{x})=\alpha$ 0.672 0.932 0.575 0.710 0.304
No site comp. $k_{H}=1$ 0.651 0.906 0.557 0.669 0.295
No mut. prob. $k_{p}=1$ 0.651 0.923 0.584 0.682 0.327
No mut. prob./site comp. $k_{p}=k_{H}=1$ 0.672 0.933 0.586 0.688 0.343
No inter-residue dist. $k_{d}=1$ 0.696 0.965 0.599 0.702 0.301
No sequence kernel $k_{\text{seq}}=0$ 0.673 0.926 0.622 0.718 0.440
No structure kernel $k_{\text{struct}}=0$ 0.712 0.993 0.626 0.734 0.358
[…]
Kermut (SE in $k_{H}$) $k_{H}=k_{\text{SE}}$ 0.648 0.899 0.558 0.669 0.299
Kermut (JSD in $k_{H}$) $k_{H}=k_{\text{JSD}}$ 0.642 0.893 0.558 0.667 0.291
Kermut (product) $k=k_{\text{struct}}\times k_{\text{seq}}$ 0.643 0.884 0.548 0.667 0.283
Kermut 0.640 0.903 0.558 0.665 0.289
Appendix G Results for multi-mutants in ProteinGym
[…]
Table G.1: Results in multi-mutant setting. Each row corresponds to a different setting of training and evaluation domain. Third row corresponds to the fold_rand_multiples split-scheme from ProteinGym. Experiments are carried out on 51 datasets, corresponding to all datasets with multiple mutants with less than 7500 variants in total with the exception of GCN4_YEAST_Staller_2018, which has been removed due to its high mutation count. : Constant mean.
@@ Appendix G Results for multi-mutants in ProteinGym
Domain Kermut Kermut $k_{\text{seq}}$ Kermut Kermut $k_{\text{seq}}$
@@
Appendix H Results on GB1 landscape from the FLIP benchmark
To further investigate Kermut’s performance in a multi-mutant setting, we apply it to the GB1 fitness landscape from FLIP [6]. The results can be seen in Table H.1. Kermut’s base configuration severely underperforms in the 1-vs-rest split, while reaching similar correlation coefficients in the 2-vs-rest and 3-vs-rest splits. A reason for the initial low score might be the that the accuracy of zero-shot methods at different mutation orders tend to decrease with higher mutation count. Despite this, we still see low performance compared to other models when removing the zero-shot mean function.
The GB1 landscape is comprised of 149,361 mutations at exactly four highly epistatic positions in the GB1 binding domain of Protein G. This is a challenging task for Kermut, whose structural kernel directly compares sites. With only four sites to compare, Kermut fails to accurately model the fitness landscape. Additionally, as described in the main text, the only epistatic modeling in Kermut is via the mean-pooled protein language model embeddings in the sequence kernel, which proves insufficient to capture the interplay of these highly epistatic sites. Further experimentation is required to thoroughly gauge Kermut’s performance across diverse multi-mutant assays where the number of mutated residues is variable and epistatis plays a central role.
Table H.1: Performance on FLIP’s GB1 landscape. : Reference results from FLIP. Best and second best scores per split has been highlighted.
Model 1-vs-rest 2-vs-rest 3-vs-rest low-vs-high
ESM-1b (per AA) 0.28 0.55 0.79 0.59
ESM-1b (mean) 0.32 0.36 0.54 0.13
ESM-1b (mut mean) -0.08 0.19 0.49 0.45
ESM-1v (per AA) 0.28 0.28 0.82 0.51
ESM-1v (mean) 0.32 0.32 0.77 0.10
ESM-1v (mut mean) 0.19 0.19 0.80 0.49
ESM-untrained (per AA) 0.06 0.06 0.48 0.23
ESM-untrained (mean) 0.05 0.05 0.46 0.10
ESM-untrained (mut mean) 0.21 0.21 0.57 0.13
Ridge 0.28 0.59 0.76 0.34
CNN 0.17 0.32 0.83 0.51
Levenshtein 0.17 0.16 -0.04 -0.10
BLOSUM62 0.15 0.14 0.01 -0.13
Kermut -0.14 0.52 0.77 0.35
Kermut (constant mean) 0.37 0.55 0.77 0.36
Baseline GP 0.40 0.57 0.73 0.42
Appendix I Histogram over assays for multi-mutant datasets
A histogram over normalized assay values for 51/69 dataset with multi-mutants (total fewer than 7500 sequences) can be seen in Figure I.1. The histograms are colored according to the number of mutations per variant. The assay distribution belong to different modalities depending on the number of mutations present, where double mutations often lead to a loss of fitness.
Figure I.1: Histogram over normalized assay values for 51/69 datasets with multi-mutants. All datasets with more than 7500 variants are ignored. The histograms are colored according to the number of mutations per variant. The assay distribution belong to different modalities depending on the number of mutations present, where double mutations commonly lead to a loss of fitness.
Appendix J Uncertainty calibration
J.1 Confidence interval-based calibration
@@ Appendix I Histogram over assays for multi-mutant datasets
J.2 Error-based calibration
@@
J.3 Uncertainty calibration across all datasets
@@
Figure J.1: Calibration metrics per domain for Kermut and the sequence kernel on ESM-2 embeddings. Random, modulo, and contiguous domains are from the ProteinGym substitution benchmark. Multiples corresponds to training and testing on both single and double mutants. Extrapolation corresponds to training on singles and predicting on doubles. 51 datasets with multi-mutants was used for the figure for all domains for comparability. The performance results for the multi-mutant setting can be found in Table G.1. Errorbars correspond to standard error.
@@
For each dataset, split, and fold, we perform a linear regression to the error-based calibration curves and summarize the slope and intercept. As perfect calibration corresponds to a diagonal line, we want the distribution over the slopes to be centred on one and the distribution over intercepts to be centred on zero. Boxplots for this analysis can be seen in Figure J.2. As indicated in (a), most of Kermut’s calibration curves are well-behaved, while the baseline GP in (b) is generally poorly calibrated.
@@
Figure J.2: Boxplot of intercepts and slopes of error-based calibration curves for Kermut and a baseline GP with the sequence kernel on ESM-2 embeddings. Perfect calibration has an intercept of zero and a slope of one (indicated by dashed lines). The baseline GP has poor calibration compared to Kermut. Horizontal lines indicate 0.9, 0.75, 0.5, 0.25, 0.1 quantiles.
Appendix K Predicted vs. true values for calibration datasets
K.1 BLAT_ECOLX_Stiffler_2015
Figure K.1: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
K.2 PA_I34A1_Wu_2015
Figure K.2: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
K.3 TCRG1_MOUSE_Tsuboyama_2023
Figure K.3: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
K.4 OPSD_HUMAN_Wan_2019
Figure K.4: Predicted means (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
Appendix L Calibration curves for ProteinNPT
L.1 ProteinNPT details
@@
Figure L.1: Calibration curves for ProteinNPT on the four dataset from Table 3. Standard deviation over CV folds is shown as vertical bars. Perfect calibration corresponds to diagonal lines ($y=x$) and is shown as dashed lines in each plot. The predictive uncertainties for ProteinNPT are very small, resulting in poor out-of-the-box calibration as seen on the x-axis in (b) and in Figures L.2, L.3, L.5 and L.4. However, as indicated by the trend in both (a) and (b), the errors correlate with the magnitude of the uncertainties. This suggests that a recalibration might be sufficient to achieve good calibration.
L.3 Predicted vs. true values for calibration datasets
L.3.1 BLAT_ECOLX_Stiffler_2015
Figure L.2: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$). While the predictions are good, the model is very overconfident.
L.3.2 PA_I34A1_Wu_2015
Figure L.3: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$). Despite the relatively poor predictions, the model remains overconfident.
L.3.3 TCRG1_MOUSE_Tsuboyama_2023
Figure L.4: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$). While the predictions are good, the model is very overconfident. Despite the relatively poor predictions, the model remains overconfident.
L.3.4 OPSD_HUMAN_Wan_2019
Figure L.5: Predicted means by ProteinNPT (±$\pm 2\sigma$) vs. true values. Columns correspond to CV-schemes. Rows correspond to test folds. Perfect prediction corresponds to dashed diagonal line ($x=y$).
Appendix M Alternative zero-shot methods
@@
Table M.1: Performance using alternate zero-shot methods. The experiments are carried out on 174/217 datasets. : ESM-2 in this table is equivalent to Kermut from the main results in Table 1.
@@
EVE 0.608 0.627 0.750 0.662 0.731 0.682 0.412 0.608
ESM-2 0.605 0.628 0.743 0.659 0.730 0.683 0.420 0.611
GEMME 0.605 0.622 0.744 0.657 0.728 0.682 0.416 0.609
VESPA 0.606 0.623 0.742 0.657 0.737 0.698 0.424 0.620
TranceptEVE L 0.600 0.619 0.744 0.654 0.741 0.693 0.420 0.618
ESM-IF 0.583 0.606 0.738 0.642 0.757 0.708 0.424 0.630
ProteinMPNN 0.575 0.599 0.734 0.636 0.769 0.718 0.429 0.639
Constant mean 0.569 0.596 0.735 0.633 0.770 0.718 0.429 0.639
Appendix N Hyperparameter visualization
N.1 Hyperparameter distributions
The distributions of Kermut’s hyperparameters across a number of datasets from ProteinGym can be seen divided by split scheme in Figure N.1. $\lambda$ is a scale parameter for the structure kernel, while $\pi$ is a balancing parameter which lets the model assign importance to the structure and sequence kernels, respectively. $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are scale coefficients in the kernels’ exponents. Their inverses are shown to facilitate easier comparison with the squared exponential kernel’s lengthscale, $l_{\text{SE}}$.
Figure N.1: Distributions of Kermut’s hyperparameter across ProteinGym assays and splits. The inverses of $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are shown to facilitate easier comparison with the sequence kernel’s lengthscale.
N.2 Hyperparameters vs. dataset size
Figure N.2: Distributions of Kermut’s hyperparameter across ProteinGym assays and splits visualized against dataset sizes. The inverses of $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are shown to facilitate easier comparison with the sequence kernel’s lengthscale.
N.3 Hyperparameters vs. sequence length
Figure N.3: Distributions of Kermut’s hyperparameter across ProteinGym assays and splits visualized against sequence length. The inverses of $\gamma_{1}$, $\gamma_{2}$, and $\gamma_{3}$ are shown to facilitate easier comparison with the sequence kernel’s lengthscale.
N.4 $\pi$ vs. $\lambda$
Figure N.4: The structure kernel scale hyperparameter, $\lambda$, is shown against the kernel balancing hyperparameter, $\pi$.
@@
We show the performance of Kermut and ProteinNPT per DMS assay in Figures O.1, O.2, O.3 and O.4. The figures show the average performance and the performance per split, respectively.
@@
Figure O.1: Spearman’s correlation per DMS assay, averaged over the three split schemes. The figure has been split in two to fit.
@@
Figure O.2: Spearman’s correlation per DMS assay in the random split scheme. The figure has been split in two to fit. The performance difference between Kermut and ProteinNPT for the random split is smallest (a), i.e., for the assays where the performance is high.